OpenStudy (anonymous):

help!!Solve. x2 + 5x + 6 = 0

OpenStudy (anonymous):

I got -2,-3

OpenStudy (anonymous):

I got the same.

OpenStudy (anonymous):

That's correct. It's -2 and -3. (x-(-3))(x-(-2)) = (x+3)(x+2) = x² + 5x + 6